OpenStudy (anonymous):

Help Please =D Which of the following equations has an infinite number of solutions? A) 3x – 3 = –4x B) 2y + 4 – y = 16 C) 7x + 5 = 4x + 5 + 3x D) 6y – 2 = 2(y – 1)

OpenStudy (anonymous):

7x + 5 = 4x + 5 + 3x 7x+5=7x+5

OpenStudy (anonymous):

So, what do u think?

OpenStudy (anonymous):

um :p

OpenStudy (anonymous):

For any value of x 7x+5=7x+5 this is TRUE

OpenStudy (anonymous):

Else u can pick any value two value for x and y and check which equations are satisfied for both values